Solve the differential equation: Y + y = sec(x) given that the complimentary solution is ye(s) = ci cos(x) + c2 sin(x)
Answer #1

Solving the differential equation we getThe given differential equation is y+y = Seeln) Let y = Aemn (A&O) be the trial solution of the equation 2. y= du a au (AemCoslr) sin(n) 11 -sinin Costw = cos(n) + sineen) = 10 ^y, and y, are linearly independent. Let yp = 9, (n) (4) +90) U₂ (1) Whi. Yp=4, (4) 4 (4) + 4,142 (4) * Yp = cos(n). Ini cos(WI+ sining.n. . The general solution is y = y + Yp y G COS(M) tesinlus